# Adventures in Algebraic Geometry

### v1

@@
It can be shown that all del Pezzo surfaces are either ${\mathbb{C}}{\mathbb{P}}^{1}\times{\mathbb{C}}{\mathbb{P}}^{1}$ or blowups of at most $8$ distinct points $p_{1},\ldots,p_{k}$ on ${\mathbb{C}}{\mathbb{P}}^{2}$, provided that these points are in a sufficiently general position (for example, three points can not be collinear). An interesting feature of a del Pezzo surface is its set of $(-1)$-curves, which are the smooth curves of genus zero that can be blown down to get another del Pezzo surface. We will comment on these below.
@@
  • $k=1.$ Blowing up one point on ${\mathbb{C}}{\mathbb{P}}^{2}$, we get the Hirzebruch surface $H_{1}$. It has a unique $(-1)$-curve, namely the exceptional curve of the blowdown morphism.
@@
  • $k=2.$ When we blow up two points $p_{1}$ and $p_{2}$ on ${\mathbb{C}}{\mathbb{P}}^{2}$, we get three $(-1)$-curves: two exceptional curves $E_{1}$ and $E_{2}$ that blow down to $p_{1}$ and $p_{2}$ respectively, and the proper preimage $l$ of the line $\overline{p_{1}p_{2}}\subset{\mathbb{C}}{\mathbb{P}}^{2}$. We can blow down $l$, and the resulting surface turns out to be isomorphic to ${\mathbb{C}}{\mathbb{P}}^{1}\times{\mathbb{C}}{\mathbb{P}}^{1}$. The reader can try to prove it now, or wait until we talk about toric varieties in Section 29.
@@
  • $k=3.$ Since points can not be collinear, we might as well choose to blow up $(1:0:0)$, $(0:1:0)$ and $(0:0:1)$. There are now six $(-1)$-curves in $X$, connected to each other like vertices in a hexagon. They are three exceptional curves and three proper preimages of the coordinate lines. In fact, we have seen this surface in Exercise 2 of Section 15.
@@
  • $k=4.$ There is only one such surface, since every four points on ${\mathbb{C}}{\mathbb{P}}^{2}$, no three of which are collinear, can be moved to $(1:0:0)$, $(0:1:0)$, $(0:0:1)$ and $(1:1:1)$ by the action of ${\rm PGL}(3,{\mathbb{C}})$. This surfaces has ten $(-1)$-curves: $4$ exceptional curves and $6$ proper preimages of the lines through two out of $4$ points.
@@
  • $k=5.$ There is now a two-dimensional family of surfaces, which we looked at in Exercise 3 of Section 19. There are $16$ $(-1)$-curves: $5$ exceptional curves of the blowup, $10$ proper preimages of lines and one proper preimage of the conic through $p_{1},\ldots,p_{5}$. Sections of the anticanonical line bundle on $X$ can be identified with cubic polynomials in the homogeneous coordinates of ${\mathbb{C}}{\mathbb{P}}^{2}$ which are zero on $p_{1},\ldots,p_{5}$. This is a five-dimensional space, and it gives an embedding of $X$ into ${\mathbb{C}}{\mathbb{P}}^{4}$, so that the ideal of $X$ is generated by two degree two equations.
@@
  • $k=6.$ This case is a bit more famous than the others. There are $27$ $(-1)$-curves of $X$, which come from $6$ exceptional curves, $15$ proper preimages of lines and $6$ proper preimages of conics through $5$ out of six points we blow up. All such surfaces are embedded into ${\mathbb{C}}{\mathbb{P}}^{3}$ by the sections of $K_{X}^{\vee}$ and the image is given by one cubic equation. The $(-1)$-curves become lines in ${\mathbb{C}}{\mathbb{P}}^{3}$ in this embedding. Conversely, every smooth cubic surface in ${\mathbb{C}}{\mathbb{P}}^{3}$ is del Pezzo, has $27$ lines on it, and any collection of pairwise disjoint lines on it can be contracted to get ${\mathbb{C}}{\mathbb{P}}^{2}$. We do not prove these classical results, but we will see how the expected number of lines can be deduced using calculations in the cohomology ring of the Grassmannian of lines in Exercise 3 of Section 26.
@@
  • $k=7.$ In this case, the surface $X$ always has an involution, so that the quotient is ${\mathbb{C}}{\mathbb{P}}^{2}$ and the map $X\to{\mathbb{C}}{\mathbb{P}}^{2}$ is ramified over a smooth curve of degree $4$. There are $56$ $(-1)$-curves on $X$. In addition to the $7$ exceptional curves, proper preimages of $21$ lines through two of the blowup points and $21$ conics through five of the points, there are also proper preimages of $7$ cubic curves in ${\mathbb{C}}{\mathbb{P}}^{2}$ which pass through $6$ out of seven points of the blowup and are singular at the $7$-th point. The anticanonical line bundle is not very ample, i.e. sections of $K_{X}^{\vee}$ do not define an embedding.
@@
  • $k=8.$ There are $240$ $(-1)$-curves. Specifically, there are $8$ exceptional curves, $28$ and $56$ proper preimages of lines and conics respectively and $64$ proper preimages of cubic curves which are singular at one of the points and pass through six points of the blowup. Furthermore, there are proper preimages of $56$ degree four curves which are singular at three points and pass through the remaining five points, and $28$ degree five curves which are singular at six points and pass through the remaining two. The anticanonical line bundle is not very ample.
@@
Exercise 3. A $(-1)$-curve in a smooth surface $X$ is, by definition, isomorphic to ${\mathbb{C}}{\mathbb{P}}^{1}$ and satisfies $E^{2}=-1$. Use adjunction formula to prove that it satisfies $K\cdot E=-1$ where $K$ is the Weil divisor that corresponds to the canonical line bundle $K_{X}\to X$.
@@
is not a birational invariant. The proof that works for $m>0$ is not applicable, because vector fields do not pull back. For instance, for $m=-1$,
@@
After the dimension, it is arguably the most important invariant of an algebraic variety. Occasionally, $\kappa(X)=-\infty$ is denoted by $\kappa(X)=-1$, but the $\infty$- ∞ notation is preferable, because it satisfies
@@
  • Consider the double cover $X\to{\mathbb{C}}{\mathbb{P}}^{2}$ ramified over a smooth degree $6$ curve $f(x_{0},x_{1},x_{2})=0$ which roughly means looking at $t^{2}=f(x_{0},x_{1},x_{2})$ in some appropriate space, see Exercise 2.21.421.421.4 We do not explain why $K_{X}\to X$ is trivial in this case. The parameter count goes as follows. We have a $28$-dimensional space of homogeneous polynomials of degree $6$, which gives $27$ parameters up to scaling. Moreover, we should subtract the dimension of the automorphism group of ${\mathbb{C}}{\mathbb{P}}^{2}$, which is $\dim{\rm PGL}(3,{\mathbb{C}})=8$. Thus, we get $27-8=19$.
@@
and call $x_{i}$ the Chern roots of $X$22.222.222.2 I have also seen these more properly referred to as the Chern roots of the tangent bundle to $X$. Perhaps this will become the common convention in the future.
@@
We can think of vector fields on ${\mathbb{C}}{\mathbb{P}}^{n}$ as ${\mathbb{C}}^{*}$-equivariant vector fields on ${\mathbb{C}}^{n+1}\setminus\{{\mathbf{0}}\}$ of “weight $1$”, up to the vector field $\sum_{i=0}^{n}x_{i}\frac{\partial}{\partial x_{i}}$ along the fibers of the quotient map. ∎
@@
$0\to\mathcal{O}\to\mathcal{O}(1)^{\oplus n+1}\to TX\to 0$
@@
$0\to(TX)^{\vee}\to\mathcal{O}(-1)^{\oplus n+1}\to\mathcal{O}\to 0$
The line bundle $\mathcal{O}$ is the trivial bundle ×${\mathbb{C}}\times X\to X$, whose global sections are the holomorphic functions on $X$, which are constants. So $\Gamma(X,\mathcal{O})\cong{\mathbb{C}}$. We claim that the line bundle $\mathcal{O}(-1)\to{\mathbb{C}}{\mathbb{P}}^{n}$ has only zero holomorphic sections. One way of seeing it is that the coordinates $x_{0},\ldots,x_{n}$ on ${\mathbb{C}}{\mathbb{P}}^{n}$ are sections of $\mathcal{O}(1)$, so for any section $s\in\Gamma({\mathbb{C}}{\mathbb{P}}^{n},\mathcal{O}(-1))$ the tensor product $s\otimes x_{0}$ is a section of $\mathcal{O}$ and is therefore a constant. Since it is zero at $x_{0}=0$, we see that it must be a zero constant. This is implies that $s$ is zero for $x_{0}\neq 0$ and is thus zero everywhere by continuity.
@@
[…]
@@
    Here we used Proposition 14.6 that the degree of the canonical line bundle is $22g-$ 2, and the fact that $TX$ is the dual of the canonical line bundle.
@@
[…]
@@
Exercise 2. Use Kodaira vanishing theorem to show that for all $k>-n-1$ we have $H^{>0}({\mathbb{C}}{\mathbb{P}}^{n},\mathcal{O}(k))=0$. Hint: Use Exercise 3 of Section 9.
@@
Clearly, for $k=0$ or $k=n$ the set ${\rm Gr}(k,V)$ consists of one element. For $k=1$ we have ${\rm Gr}(k,V)={\mathbb{P}}V\cong{\mathbb{C}}{\mathbb{P}}^{n-1}$. For $k=n-1$ we have ${\rm Gr}(k,V)={\mathbb{P}}V^{\vee}\cong{\mathbb{C}}{\mathbb{P}}^{n-1}$. This makes ${\rm Gr}(2,4)$ the first interesting case.
@@
It can be seen from the same construction that the map is an embedding. If two points go to the same point in the projective space, then they must have the same nonzero minor, which we can for simplicity assume to be $M_{1,\ldots,k}$. We can assume that $M$ is in the reduced form and observe that determinants of minors that use $(k-1)$ of the first $k$ columns pick up precisely the coordinates of the reduced form of $M$. It is also easy to see that tangent vectors do not collapse under $\mu$.
@@
$\begin{array}[]{l}{\rm contr}_{\alpha}(v_{1}\wedge\cdots\wedge v_{m})=\alpha(v_{1})v_{2}\wedge\cdots\wedge v_{m}\\ \hskip 20.0pt-\alpha(v_{2})v_{1}\wedge v_{3}\wedge\cdots\wedge v_{m}+\ldots+(-1)^{m-1}\alpha(v_{m})v_{1}\wedge\cdots\wedge v_{m-1}.\end{array}$ (24.1)
@@
$l\subset\{(0:*:*:*)\}\hskip 10.0pt(0:0:0:1)\in l\subset\{(0:*:*:*)\}\hskip 10.0ptl=(0:0:*:*)$ (25.15)
@@
which allows us to compute $q_{i}$ in terms of $s_{1}$ and $s_{2}$. It also gives us relations on $s_{i}$ by looking at the coefficients of degree $4$, $5$ and $6$.
@@
$\begin{array}[]{c}1\\ 0\hskip 24.0pt0\\ 0\hskip 24.0pt1\hskip 24.0pt0\\ 1\hskip 17.0pt101\hskip 13.0pt101\hskip 17.0pt1\\ 0\hskip 24.0pt1\hskip 24.0pt0\\ 0\hskip 24.0pt0\\ 1\end{array}$
@@
$\begin{array}[]{c}1\\ 0\hskip 24.0pt0\\ 0\hskip 20.0pt101\hskip 20.0pt0\\ 1\hskip 24.0pt1\hskip 24.0pt1\hskip 24.0pt1\\ 0\hskip 20.0pt101\hskip 20.0pt0\\ 0\hskip 24.0pt0\\ 1\end{array}$
@@
It is common to put a condition $\mathcal{F}(\emptyset)=0$ for both presheaves and sheaves. Technically, since we can do an empty cover of an empty set, the uniqueness of gluing implies this statement for sheaves.27.127.127.1 This reminds me of the statement that the determinant of a $0\times 0$ matrix is $1$, because it is the sum of $0!=1$ terms, the term is a product of 00 entries (so it is equal to $1$), and the sign of the term is positive because there are zero inversions.
@@
which sends the prime ideal $p\subset B$ to the prime ideal $f^{-1}(p)\subset A$27.327.327.3 Preimages of prime ideals are prime, while the same can not be said about the maximal ideals. This is a good reason why we want to include all prime ideals into ${\rm Spec}$. This is a continuous map of topological spaces, see Exercise 2. There is also some additional data that mimics the pullback of holomorphic functions for holomorphic maps of complex manifolds.
@@
For any oriented basis $(l_{1},l_{2})$ of $L$, we define $\tau=\frac{l_{2}}{l_{1}}$ in the upper half plane =$\mathcal{H}=\{{\rm Im}\,\tau>0\}$ and observe that
@@
${\mathbb{R}}$$|\tau|=1$${\mathrm{i}}$${\mathrm{e}}^{2\pi{\mathrm{i}}/3}$$\mathcal{H}$00
We do not prove that this is the fundamental domain but instead refer to Chapter 7 of [27]. Most points of $\mathcal{H}$ have the stabilizer {±$\{\pm{\rm Id}\}$ which corresponds to the Kummer involution $z\mapsto(-z)$ on ${\mathbb{C}}/L$. For $\tau={\mathrm{i}}$, the stabilizer is isomorphic to ${\mathbb{Z}}/4{\mathbb{Z}}$, generated by $\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right)$. This corresponds to the square lattice of Gaussian integers with the extra automorphism that is the multiplication by ${\mathrm{i}}$. Of course, we have conjugate stabilizers in the ${\rm SL}(2,{\mathbb{Z}})$ orbit of $\tau={\mathrm{i}}$. Similarly, for $\tau={\mathrm{e}}^{\frac{2\pi{\mathrm{i}}}{3}}$, the stabilizer is ${\mathbb{Z}}/6{\mathbb{Z}}$, generated by $\left(\begin{array}[]{rr}1&1\\ -1&0\end{array}\right)$. The lattice is the hexagonal lattice of Eisenstein integers, again with more automorphisms.
@@
We remark that the weight of a nonzero modular form for ${\rm SL}(2,{\mathbb{Z}})$ has to be even, because $-{\rm id}\in{\rm SL}(2,{\mathbb{Z}})$ implies that $f(\tau)=(-1)^{k}f(\tau)$.
@@
which equals $(c\tau+d)^{k}G_{k}(\tau)$ by reindexing the summation. The first statement then follows from modularity (or just observing that $(m,n)$ and $(-m,-n)$ terms cancel out). ∎
@@
respectively. One can, and does, talk about modular forms of weight $k$ with respect to these groups. The quotients of $\mathcal{H}$ by these groups are compactified by adding a finite number of cusps, corresponding to the orbits of the group on ${\mathbb{Q}}\sqcup\{{\mathrm{i}}\infty\}$. The resulting curves $X(n)$, $X_{1}(n)$ and $X_{0}(n)$ have genus $g$ that grows roughly as $c\,n^{3}$, $c\,n^{2}$ and $c\,n$ respectively.
@@
[…]
@@
The set $\sigma\cap M$ has a natural semigroup structure. We denote by ${\mathbb{C}}[\sigma\cap M]$ the corresponding semigroup ring. It is a vector space with a basis indexed by $m\in\sigma\cap M$. We denote the basis elements by $[m]$ and define the multiplication by $[m_{1}][m_{2}]=[m_{1}+m_{2}]$, naturally extended to the whole semigroup ring.
@@
  • More generally, if $\sigma={\mathbb{R}}_{\geq 0}(1,0)+{\mathbb{R}}_{\geq 0}(n-1,n)$, then ${\mathbb{C}}[\sigma\cap M]$ is generated by $u=[(1,0)]$, $v=[(n-1,n)]$, $w=[(1,1)]$ which satisfy the relation $uv-w^{n}=0,$ and we get

### v2

@@
It can be shown that all del Pezzo surfaces are either ${\mathbb{C}}{\mathbb{P}}^{1}\times{\mathbb{C}}{\mathbb{P}}^{1}$ or blowups of at most $8$ distinct points $p_{1},\ldots,p_{k}$ on ${\mathbb{C}}{\mathbb{P}}^{2}$, provided that these points are in a sufficiently general position (for example, three points can not be collinear). An interesting feature of a del Pezzo surface is its set of (1)(-1)-curves, which are the smooth curves of genus zero that can be blown down to get another del Pezzo surface. We will comment on these below.
@@
  • $k=1.$ Blowing up one point on ${\mathbb{C}}{\mathbb{P}}^{2}$, we get the Hirzebruch surface $H_{1}$. It has a unique (1)(-1)-curve, namely the exceptional curve of the blowdown morphism.
@@
  • $k=2.$ When we blow up two points $p_{1}$ and $p_{2}$ on ${\mathbb{C}}{\mathbb{P}}^{2}$, we get three (1)(-1)-curves: two exceptional curves $E_{1}$ and $E_{2}$ that blow down to $p_{1}$ and $p_{2}$ respectively, and the proper preimage $l$ of the line $\overline{p_{1}p_{2}}\subset{\mathbb{C}}{\mathbb{P}}^{2}$. We can blow down $l$, and the resulting surface turns out to be isomorphic to ${\mathbb{C}}{\mathbb{P}}^{1}\times{\mathbb{C}}{\mathbb{P}}^{1}$. The reader can try to prove it now, or wait until we talk about toric varieties in Section 29.
@@
  • $k=3.$ Since points can not be collinear, we might as well choose to blow up $(1:0:0)$, $(0:1:0)$ and $(0:0:1)$. There are now six (1)(-1)-curves in $X$, connected to each other like vertices in a hexagon. They are three exceptional curves and three proper preimages of the coordinate lines. In fact, we have seen this surface in Exercise 2 of Section 15.
@@
  • $k=4.$ There is only one such surface, since every four points on ${\mathbb{C}}{\mathbb{P}}^{2}$, no three of which are collinear, can be moved to $(1:0:0)$, $(0:1:0)$, $(0:0:1)$ and $(1:1:1)$ by the action of ${\rm PGL}(3,{\mathbb{C}})$. This surfaces has ten (1)(-1)-curves: 44 exceptional curves and $6$ proper preimages of the lines through two out of $4$ points.
@@
  • $k=5.$ There is now a two-dimensional family of surfaces, which we looked at in Exercise 3 of Section 19. There are $16$ (1)(-1)-curves: 55 exceptional curves of the blowup, $10$ proper preimages of lines and one proper preimage of the conic through $p_{1},\ldots,p_{5}$. Sections of the anticanonical line bundle on $X$ can be identified with cubic polynomials in the homogeneous coordinates of ${\mathbb{C}}{\mathbb{P}}^{2}$ which are zero on $p_{1},\ldots,p_{5}$. This is a five-dimensional space, and it gives an embedding of $X$ into ${\mathbb{C}}{\mathbb{P}}^{4}$, so that the ideal of $X$ is generated by two degree two equations.
@@
  • $k=6.$ This case is a bit more famous than the others. There are $27$ (1)(-1)-curves of $X$, which come from $6$ exceptional curves, $15$ proper preimages of lines and $6$ proper preimages of conics through $5$ out of six points we blow up. All such surfaces are embedded into ${\mathbb{C}}{\mathbb{P}}^{3}$ by the sections of $K_{X}^{\vee}$ and the image is given by one cubic equation. The (1)(-1)-curves become lines in ${\mathbb{C}}{\mathbb{P}}^{3}$ in this embedding. Conversely, every smooth cubic surface in ${\mathbb{C}}{\mathbb{P}}^{3}$ is del Pezzo, has $27$ lines on it, and any collection of pairwise disjoint lines on it can be contracted to get ${\mathbb{C}}{\mathbb{P}}^{2}$. We do not prove these classical results, but we will see how the expected number of lines can be deduced using calculations in the cohomology ring of the Grassmannian of lines in Exercise 3 of Section 26.
@@
  • $k=7.$ In this case, the surface $X$ always has an involution, so that the quotient is ${\mathbb{C}}{\mathbb{P}}^{2}$ and the map $X\to{\mathbb{C}}{\mathbb{P}}^{2}$ is ramified over a smooth curve of degree $4$. There are $56$ (1)(-1)-curves on $X$. In addition to the $7$ exceptional curves, proper preimages of $21$ lines through two of the blowup points and $21$ conics through five of the points, there are also proper preimages of $7$ cubic curves in ${\mathbb{C}}{\mathbb{P}}^{2}$ which pass through $6$ out of seven points of the blowup and are singular at the $7$-th point. The anticanonical line bundle is not very ample, i.e. sections of $K_{X}^{\vee}$ do not define an embedding.
@@
  • $k=8.$ There are $240$ (1)(-1)-curves. Specifically, there are $8$ exceptional curves, $28$ and $56$ proper preimages of lines and conics respectively and $64$ proper preimages of cubic curves which are singular at one of the points and pass through six points of the blowup. Furthermore, there are proper preimages of $56$ degree four curves which are singular at three points and pass through the remaining five points, and $28$ degree five curves which are singular at six points and pass through the remaining two. The anticanonical line bundle is not very ample.
@@
Exercise 3. A (1)(-1)-curve in a smooth surface $X$ is, by definition, isomorphic to ${\mathbb{C}}{\mathbb{P}}^{1}$ and satisfies $E^{2}=-1$. Use adjunction formula to prove that it satisfies $K\cdot E=-1$ where $K$ is the Weil divisor that corresponds to the canonical line bundle $K_{X}\to X$.
@@
is not a birational invariant. The proof that works for $m>0$ is not applicable, because vector fields do not pull back. For instance, for m=1m=-1,
@@
After the dimension, it is arguably the most important invariant of an algebraic variety. Occasionally, $\kappa(X)=$-$\infty$ is denoted by $\kappa(X)=-1$, but the -$\infty$ notation is preferable, because it satisfies
@@
  • Consider the double cover $X\to{\mathbb{C}}{\mathbb{P}}^{2}$ ramified over a smooth degree $6$ curve $f(x_{0},x_{1},x_{2})=0$ which roughly means looking at $t^{2}=f(x_{0},x_{1},x_{2})$ in some appropriate space, see Exercise 2.21.421.421.4We do not explain why $K_{X}\to X$ is trivial in this case. The parameter count goes as follows. We have a $28$-dimensional space of homogeneous polynomials of degree $6$, which gives $27$ parameters up to scaling. Moreover, we should subtract the dimension of the automorphism group of ${\mathbb{C}}{\mathbb{P}}^{2}$, which is $\dim{\rm PGL}(3,{\mathbb{C}})=8$. Thus, we get 278=1927-8=19.
@@
and call $x_{i}$ the Chern roots of XX.22.222.222.2I have also seen these more properly referred to as the Chern roots of the tangent bundle to $X$. Perhaps this will become the common convention in the future.
@@
[…]
@@
$0\to\mathcal{O}\to\mathcal{O}(1)^{\oplus n+1}\to TX\to 0.$
@@
$0\to(TX)^{\vee}\to\mathcal{O}(-1)^{\oplus n+1}\to\mathcal{O}\to 0.$
The line bundle $\mathcal{O}$ is the trivial bundle ${\mathbb{C}}\times X\to X$, whose global sections are the holomorphic functions on $X$, which are constants. So $\Gamma(X,\mathcal{O})\cong{\mathbb{C}}$. We claim that the line bundle $\mathcal{O}(-1)\to{\mathbb{C}}{\mathbb{P}}^{n}$ has only zero holomorphic sections. One way of seeing it is that the coordinates $x_{0},\ldots,x_{n}$ on ${\mathbb{C}}{\mathbb{P}}^{n}$ are sections of $\mathcal{O}(1)$, so for any section $s\in\Gamma({\mathbb{C}}{\mathbb{P}}^{n},\mathcal{O}(-1))$ the tensor product $s\otimes x_{0}$ is a section of $\mathcal{O}$ and is therefore a constant. Since it is zero at $x_{0}=0$, we see that it must be a zero constant. This is implies that $s$ is zero for $x_{0}\neq 0$ and is thus zero everywhere by continuity.
@@
$\begin{array}[]{l}0\to\Gamma(X,W_{1})\to\Gamma(X,W_{2})\to\Gamma(X,W_{3})\to\\[5.0pt] \to H^{1}(X,W_{1})\to H^{1}(X,W_{2})\to H^{1}(X,W_{3})\to\\[5.0pt] \to H^{2}(X,W_{1})\to\cdots\end{array}$ (23.1)
@@
    Here we used Proposition 14.6 that the degree of the canonical line bundle is 2g22g-2, and the fact that $TX$ is the dual of the canonical line bundle.
@@
    $={\rm Res}_{H=0}\frac{{\mathrm{e}}^{kH}}{(1-{\mathrm{e}}^{-H})^{n+1}}=\frac{1}{2\pi{\mathrm{i}}}\oint_{|H|=\varepsilon}\frac{{\mathrm{e}}^{kH}}{(1-{\mathrm{e}}^{-H})^{n+1}}\,dH$
    $\stackrel{{\scriptstyle t=1-{\mathrm{e}}^{-H}}}{{=}}\frac{1}{2\pi{\mathrm{i}}}\oint_{|t|=\varepsilon}\frac{(1-t)^{-k-1}}{t^{n+1}}\,dt=\frac{(k+1)(k+2)\cdot\ldots\cdot(k+n)}{n!}$
@@
Exercise 2. Use Kodaira vanishing theorem to show that for all k>n1k>-n-1 we have $H^{>0}({\mathbb{C}}{\mathbb{P}}^{n},\mathcal{O}(k))=0$. Hint: Use Exercise 3 of Section 9.
@@
Clearly, for $k=0$ or $k=n$ the set ${\rm Gr}(k,V)$ consists of one element. For $k=1$ we have ${\rm Gr}(k,V)={\mathbb{P}}V\cong{\mathbb{C}}{\mathbb{P}}^{n-1}$. For k=n1k=n-1 we have ${\rm Gr}(k,V)={\mathbb{P}}V^{\vee}\cong{\mathbb{C}}{\mathbb{P}}^{n-1}$. This makes ${\rm Gr}(2,4)$ the first interesting case.
@@
It can be seen from the same construction that the map is an embedding. If two points go to the same point in the projective space, then they must have the same nonzero minor, which we can for simplicity assume to be $M_{1,\ldots,k}$. We can assume that $M$ is in the reduced form and observe that determinants of minors that use (k1)(k-1) of the first $k$ columns pick up precisely the coordinates of the reduced form of $M$. It is also easy to see that tangent vectors do not collapse under $\mu$.
@@
$\begin{array}[]{l}{\rm contr}_{\alpha}(v_{1}\wedge\cdots\wedge v_{m})=\alpha(v_{1})v_{2}\wedge\cdots\wedge v_{m}\\ \qquad-\alpha(v_{2})v_{1}\wedge v_{3}\wedge\cdots\wedge v_{m}+\ldots+(-1)^{m-1}\alpha(v_{m})v_{1}\wedge\cdots\wedge v_{m-1}.\end{array}$ (24.1)
@@
$l\subset\{(0:*:*:*)\}\quad(0:0:0:1)\in l\subset\{(0:*:*:*)\}\quad l=(0:0:*:*)$ (25.15)
@@
which allows us to compute $q_{i}$ in terms of $s_{1}$ and $s_{2}$. It also gives us relations on $s_{i}$ by looking at the coefficients of degree $4$, 55 and $6$.
@@
$\begin{array}[]{c}1\\ 0\hskip 24.0pt0\\ 0\hskip 24.0pt1\hskip 24.0pt0\\ 1\qquad 101\qquad 101\qquad 1\\ 0\hskip 24.0pt1\hskip 24.0pt0\\ 0\hskip 24.0pt0\\ 1\end{array}$
@@
$\begin{array}[]{c}1\\ 0\hskip 24.0pt0\\ 0\qquad 101\qquad 0\\ 1\hskip 24.0pt1\hskip 24.0pt1\hskip 24.0pt1\\ 0\qquad 101\qquad 0\\ 0\hskip 24.0pt0\\ 1\end{array}$
@@
It is common to put a condition $\mathcal{F}(\emptyset)=0$ for both presheaves and sheaves. Technically, since we can do an empty cover of an empty set, the uniqueness of gluing implies this statement for sheaves.27.127.127.1This reminds me of the statement that the determinant of a $0\times 0$ matrix is $1$, because it is the sum of $0!=1$ terms, the term is a product of 0 entries (so it is equal to $1$), and the sign of the term is positive because there are zero inversions.
@@
which sends the prime ideal $p\subset B$ to the prime ideal $f^{-1}(p)\subset A$.27.327.327.3Preimages of prime ideals are prime, while the same can not be said about the maximal ideals. This is a good reason why we want to include all prime ideals into $\rm Spec$}. This is a continuous map of topological spaces, see Exercise 2. There is also some additional data that mimics the pullback of holomorphic functions for holomorphic maps of complex manifolds.
@@
For any oriented basis $(l_{1},l_{2})$ of $L$, we define $\tau=\frac{l_{2}}{l_{1}}$ in the upper half plane $\mathcal{H}=\{{\rm Im}\,\tau>0\}$ and observe that
@@
$\mathbb{R}$$|\tau|=1$$\mathrm{i}$${\mathrm{e}}^{2\pi{\mathrm{i}}/3}\mathcal{H}$0
We do not prove that this is the fundamental domain but instead refer to Chapter 7 of [27]. Most points of $\mathcal{H}$ have the stabilizer $\{\pm{\rm Id}\}$ which corresponds to the Kummer involution $z\mapsto(-z)$ on ${\mathbb{C}}/L$. For $\tau={\mathrm{i}}$, the stabilizer is isomorphic to ${\mathbb{Z}}/4{\mathbb{Z}}$, generated by $\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right)$. This corresponds to the square lattice of Gaussian integers with the extra automorphism that is the multiplication by $\mathrm{i}$}. Of course, we have conjugate stabilizers in the ${\rm SL}(2,{\mathbb{Z}})$ orbit of $\tau={\mathrm{i}}$. Similarly, for $\tau={\mathrm{e}}^{\frac{2\pi{\mathrm{i}}}{3}}$, the stabilizer is ${\mathbb{Z}}/6{\mathbb{Z}}$, generated by $\left(\begin{array}[]{rr}1&1\\ -1&0\end{array}\right)$. The lattice is the hexagonal lattice of Eisenstein integers, again with more automorphisms.
@@
We remark that the weight of a nonzero modular form for ${\rm SL}(2,{\mathbb{Z}})$ has to be even, because ${\rm id}\in{\rm SL}(2,{\mathbb{Z}})$ implies that $f(\tau)=(-1)^{k}f(\tau)$.
@@
which equals $(c\tau+d)^{k}G_{k}(\tau)$ by reindexing the summation. The first statement then follows from modularity (or just observing that $(m,n)$ and (m,n)(-m,-n) terms cancel out). ∎
@@
respectively. One can, and does, talk about modular forms of weight $k$ with respect to these groups. The quotients of $\mathcal{H}$ by these groups are compactified by adding a finite number of cusps, corresponding to the orbits of the group on ${\mathbb{Q}}\sqcup\{{\mathrm{i}}\infty\}$. The resulting curves $X(n)$, $X_{1}(n)$ and $X_{0}(n)$ have genus $g$ that grows roughly as $c\,n^{3}$, $c\,n^{2}$ and cnc\,n respectively.
@@
Exercise 1. Prove that if two elliptic curves ${\mathbb{C}}/L$ and ${\mathbb{C}}/L^{\prime}$ are isomorphic as complex manifolds, then there exists $\lambda\in{\mathbb{C}}$ such that $L^{\prime}=\lambda L$. Hint: Argue that the space of holomorphic one-forms on ${\mathbb{C}}/L$ is one-dimensional. Then (after arranging that the isomorphism preserves 0) compare the pullback of the differential form $dz$ on ${\mathbb{C}}/L^{\prime}$ with $dz$ on ${\mathbb{C}}/L$ and use $t=\int_{0}^{t}\,dz$.
@@
The set $\sigma\cap M$ has a natural semigroup structure. We denote by ${\mathbb{C}}[\sigma\cap M]$ the corresponding semigroup ring. It is a vector space with a basis indexed by $m\in\sigma\cap M$. We denote the basis elements by $[m]$ and define the multiplication by $m_{1}][m_{2}]=[m_{1}+m_{2}]$, naturally extended to the whole semigroup ring.
@@
  • More generally, if $\sigma={\mathbb{R}}_{\geq 0}(1,0)+{\mathbb{R}}_{\geq 0}(n-1,n)$, then ${\mathbb{C}}[\sigma\cap M]$ is generated by $u=[(1,0)]$, v=[(n1,n)]v=[(n-1,n)], $w=[(1,1)]$ which satisfy the relation $uv-w^{n}=0,$ and we get